OpenStudy (anonymous):

Yvette goes to the local bank to get $20 in quarters, dimes and nickels to use in her restaurant. She gets twice as many quarters as dimes and 10 more nickels than dimes. How many of each type of coin does she get?

OpenStudy (anonymous):

n = nickels d = dimes q = quarters .05n + .10d + .25q = 20 q = 2d n = d + 10 now sub known values into the 1st equation... .05n + .10d + .25q = 20 .05(d + 10) + .10d + .25(2d) = 20 .05d + .5 + .10d + .5d = 20 .65d + .5 = 20 .65d = 20 - .5 .65d = 19.5 d = 19.5/.65 d = 30 now sub 30 in for d q = 2d q = 2(30) q = 60 n = d + 10 n = 30 + 10 n = 40 lets check our answers... .05n + .10d + .25q = 20 .05(40) + .10(30) + .25(60) = 20 2 + 3 + 15 = 20 20 = 20 (correct) ANSWER : Yvette has 40 nickels totaling 2.00, she has 30 dimes totaling 3.00, and she has 60 quarters totaling 15.00......for a final total of $ 20.00